# Random Polyhedral Cones: Distributional Results via Gale Duality

### v1

@@ Abstract.
$\mathcal{W}_{n,d}:=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})=\{\lambda_{1}U_{1}+\ldots+\lambda_{n}U_{n}:\lambda_{1}\geq 0,\ldots,\lambda_{n}\geq 0\}.$
@@
    Let $I_{1},I_{2}\subseteq[n]$ be such that $I_{1}\cup I_{2}=[n]$ and $\max(\#I_{1},\#I_{2})\leq d-1$. We prove that the “face events” $\{\,\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I_{r})\text{ is a face of }\mathcal{W}_{n,d}\,\}$, $r=1,2$, are independent.
@@
Stochastic geometry, geometric probability, random polyhedral cones, random spherical polytopes, solid angles, Gaussian projection, Gale duality, spherical Sylvester problem, $U$-statistics, Funk–Hecke formula, spherical harmonics
@@ 1.1. Random polyhedral cones
Thus, $\mathcal{W}_{n,d}^{\circ}$ is the feasible set of a system of $n$ random linear inequalities, making it a natural and fundamental object. The random cones $\mathcal{W}_{n,d}$, $\mathcal{W}_{n,d}^{\circ}$, and related models have been studied by Cover and Efron [8], Donoho and Tanner [11], and Hug and Schneider [20]; see also [15, 21, 22, 25, 41] and Chapters 5–6 of the book by Schneider [42]. One important result is Wendel’s formula [47] (see also [43, Theorem 8.2.1]), which states that
@@ 2.1. Linear Gale duality
Gale duality (or the Gale transform) is a widely used tool in convex and discrete geometry. For detailed accounts of Gale duality and its applications, we refer to the books of Grünbaum [17, Section 5.4], Ziegler [50, Chapter 6],  Matoušek [33, Section 5.6], McMullen and Shephard [35, Chapter 3], Björner et al. [6, Section 3.4], and to the papers of McMullen [34], Shephard [46], and Grünbaum and Shephard [18]. There are two versions of Gale duality: affine and linear. The version most commonly treated in the literature is the affine one; in the present work, however, we require linear Gale duality studied in [34, 46], which we now define.
@@
    if $A\in\mathbb{R}^{d^{\prime}\times n}$ is the matrix with comumns $a_{1},\ldots,a_{n}$ and $B\in\mathbb{R}^{d^{\prime\prime}\times n}$ is the matrix with columns $b_{1},\ldots,b_{n}$, then every row of $A$ is orthogonal to every row of $B$. In matrix notation,
@@ 2.2. Gale coupling between two Gaussian projections
In this proof, we give an explicit construction of the coupling $(\mathbf{G},\mathbf{H})$. Let $X_{1},\ldots,X_{d^{\prime\prime}}$ be i.i.d. standard Gaussian points in $\mathbb{R}^{n}=\mathbb{R}^{d^{\prime}+d^{\prime\prime}}$. The linear Blaschke–Petkantschin formula [43, Theorem 7.2.1] states that for every nonnegative Borel function $h:(\mathbb{R}^{n})^{d^{\prime\prime}}\to[0,\infty)$,
@@
where $\nu_{n,d^{\prime\prime}}$ is the uniform probability distribution on $G(n,d^{\prime\prime})$, $\lambda_{L}$ is the Lebesgue measure on $L\in G(n,d^{\prime\prime})$, $\nabla_{d^{\prime\prime}}(z_{1},\ldots,z_{d^{\prime\prime}})$ denotes the $d^{\prime\prime}$-dimensional volume of the parallelepiped spanned by the vectors $z_{1},\ldots,z_{d^{\prime\prime}}$, and $B$ is a normalizing constant whose value can be found in [43, Eqn. (7.8), p. 271].
@@ 3.2. Expected face counts
As a simple consequence of Theorem 2.8, we recover the following result of Donoho and Tanner [11, Theorem 1.6]. Recall that Wendel’s formula [47] (see also [43, Theorem 8.2.1]) states that
@@
Taking up a suggestion of Gale, Schneider [41] proposed to study random polytopes $P$ obtained by choosing their (affine) Gale transform at random. (Throughout this remark, we use the term “polytope” to refer to its combinatorial type; only the combinatorial type is canonically determined by the Gale transform.) The model of Schneider [41, p. 643] depends on an even probability measure $\phi$; in the present remark we take $\phi$ to be the uniform distribution on the unit sphere in $\mathbb{R}^{k}$. We put $n=d+k$. Using the notation of Theorem 2.8, Schneider [41] considers a (d1)(d-1)-dimensional random polytope $P$ with $n$ vertices whose (affine) Gale transform is given by $\lambda_{1}V_{1},\ldots,\lambda_{n}V_{n}\in\mathbb{R}^{k}$ (with suitable $\lambda_{j}>0$) conditioned on the event $\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n})=\mathbb{R}^{k}$. By Example 2.4, this is the same conditioning event as $\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})\neq\mathbb{R}^{d}$. It then follows from Theorem 2.8 that the random cone $\mathcal{W}_{n,d}=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})$ conditioned on the event $\{\mathcal{W}_{n,d}\neq\mathbb{R}^{d}\}$, can be identified with the cone over $P$. Conversely, $P$ can be recovered as a bounded slice (cross-section) of this conditioned random cone. Such conditioned cones are called Cover–Efron cones in [20]. Consequently, the random polytopes studied in Schneider [41] have the same distribution as the bounded cross-sections of Cover–Efron cones considered in [20]. (As noted in [41, Remark on p. 648], these two objects in particular have the same expected face numbers.)
@@ 4.2. Moments of small order
In this section, we considered the cone $\mathcal{W}_{n,d}$ with $n=d$ only. Our methods do not extend to $n>d$. For example, it remains open to compute the variance of $\alpha(\mathcal{W}_{d+1,d})$.
@@ 6.1. Statement of the limit theorem
Let us record a convex-dual analogue of Theorem 6.1. Let $U_{1},\ldots,U_{n}$ be independent random vectors, each uniformly distributed on $\mathbb{S}^{d-1}$, where $n>d$. The hyperplanes $U_{1}^{\perp},\ldots,U_{n}^{\perp}$ induce a conical tessellation of $\mathbb{R}^{d}$ into polyhedral cones. The number of cones is almost surely constant and equals $C(n,d)=2\sum_{\ell=0}^{d-1}\binom{n-1}{\ell}$, by the Steiner–Schläfli formula; see [43, Lemma 8.2.1]. Choose one of these cones uniformly at random and denote it by $\mathcal{S}_{n,d}$. The random cone $\mathcal{S}_{n,d}$ is called the random Schläfli cone; see [20, Definition 3.2].
@@ 6.2. Background from the theory of $U$-statistics
The proof of Theorem 6.1 relies on the theory of $U$-statistics. In this section, we introduce the necessary notation and recall several classical results on $U$-statistics; see the books of Lee [29], Korolyuk and Borovskikh [27], and Serfling [45, Chapter 5] for further background.
@@
For $n\geq m$, the variance of $\mathbb{U}_{n}$ is given by the exact formula (see [29, Chapter 1, § 1.3, Theorem 3] or [45, Section 5.2.1, Lemma A])
@@
Equivalently, $g_{r}(x_{1},\ldots,x_{r})=g_{0}$ for all $r\in\{1,\ldots,c-1\}$ ($\mu^{r}$-a.e.), but $g_{c}(X_{1},\ldots,X_{c})$ is not a.s. constant. If $h$ is degenerate of order $c$, Equation (6.7) implies (see, e.g., [45, Section 5.3.4])
@@
The next classical theorem characterizes the weak limit of the $U$-statisic $\mathbb{U}_{n}$ under degeneracy of order $2$; see [29, Section 3.2.2], [27, Chapter 4], [40, p. 168] or [45, Section 5.5.2, p. 194].
@@
The next lemma provides an explicit formula for the conditional expectation kernel $g_{c}(v_{1},\ldots,v_{c})$ of arbitrary order. The formula is expressed in terms of conic intrinsic volumes. To each closed convex cone $C\subseteq\mathbb{R}^{k}$ one can associate conic intrinsic volumes $\upsilon_{0}(C),\ldots,\upsilon_{k}(C)$, which are nonnegative and satisfy $\upsilon_{0}(C)+\cdots+\upsilon_{k}(C)=1$. We refer to [1, 2] and [42, Section 2.3] for their definition and properties.
@@
In the conditional probability, the only random element is the random matrix $\mathbb{O}$. To compute the conditional probability, we invoke the conic kinematic formula whose proof can be found in [1, Corollary 5.2] or in [42, Theorem  4.3.5]; for applications see also [1]. The formula states that for (deterministic) polyhedral cones $A,B\subseteq\mathbb{R}^{k}$ that are not both linear subspaces,
@@
Equations (6.13) and (6.14) are proved in [15, Lemma 5.1] (building on [20, Corollary 4.3]); where it is assumed that $n\geq d$. See also [24, Theorem 5.18] for an alternative approach. We sketch a proof of (6.15) for $n\in\{1,\dots,d\}$; for background on conic intrinsic volumes and hyperplane arrangements we refer to [42, Section 5.2]. For each sign vector $\varepsilon=(\varepsilon_{1},\dots,\varepsilon_{n})\in\{\pm 1\}^{n}$, consider the random cone $C_{\varepsilon}:=\mathop{\mathrm{pos}}\nolimits(\varepsilon_{1}U_{1},\dots,\varepsilon_{n}U_{n})$. By symmetry, $\mathbb{E}\,\upsilon_{k}(C_{\varepsilon})=\mathbb{E}\,\upsilon_{k}(\mathcal{W}_{n,d})$ for every $\varepsilon$. On the other hand, with probability $1$, the cones $C_{\varepsilon}$ are precisely the chambers of a hyperplane arrangement in $\mathop{\mathrm{lin}}\nolimits(U_{1},\dots,U_{n})$ that has the same intersection poset as the coordinate hyperplane arrangement in $\mathbb{R}^{n}$, whose chambers are the $2^{n}$ orthants. Since the $k$th conic intrinsic volume of an orthant in $\mathbb{R}^{n}$ equals $\frac{1}{2^{n}}\binom{n}{k}$, the Klivans–Swartz formula [42, Theorem 5.1.4] yields $\sum_{\varepsilon\in\{\pm 1\}^{n}}\upsilon_{k}(C_{\varepsilon})=\binom{n}{k}$ a.s. Taking expectations gives $2^{n}\,\mathbb{E}\,\upsilon_{k}(\mathcal{W}_{n,d})=\sum_{\varepsilon\in\{\pm 1\}^{n}}\mathbb{E}\,\upsilon_{k}(C_{\varepsilon})=\binom{n}{k}$, which is precisely (6.15). ∎
@@
Let $U$ be uniformly distributed on $\mathbb{S}^{k-1}$. For all $v,w\in\mathbb{S}^{k-1}$ it is easy to check (see, e.g., [49, Lemma 6.7 on p. 144]) that

### v2

# Random Polyhedral Cones I: Distributional Results via Gale Duality

@@ Abstract.
$\mathcal{W}_{n,d}:=\mathop{\mathrm{pos}}(U_{1},\ldots,U_{n})=\{\lambda_{1}U_{1}+\ldots+\lambda_{n}U_{n}:\lambda_{1}\geq 0,\ldots,\lambda_{n}\geq 0\}.$
@@
    Let $I_{1},I_{2}\subseteq[n]$ be such that $I_{1}\cup I_{2}=[n]$ and $\max(\#I_{1},\#I_{2})\leq d-1$. We prove that the “face events” $\{\,\mathop{\mathrm{pos}}(U_{i}:i\in I_{r})\text{ is a face of }\mathcal{W}_{n,d}\,\}$, $r=1,2$, are independent.
@@
Stochastic geometry, geometric probability, random polyhedral cones, random spherical polytopes, solid angles, Gale duality, spherical Sylvester problem, $U$-statistics, Funk–Hecke formula, spherical harmonics
@@ 1.1. Random polyhedral cones
Thus, $\mathcal{W}_{n,d}^{\circ}$ is the feasible set of a system of $n$ random linear inequalities, making it a natural and fundamental object. The random cones $\mathcal{W}_{n,d}$, $\mathcal{W}_{n,d}^{\circ}$, and related models have been studied by Cover and Efron [8], Donoho and Tanner [11], and Hug and Schneider [20]; see also [15, 21, 22, 25, 43] and Chapters 5–6 of the book by Schneider [44]. One important result is Wendel’s formula [49] (see also [45, Theorem 8.2.1]), which states that
@@ 2.1. Linear Gale duality
Gale duality (or the Gale transform) is a widely used tool in convex and discrete geometry. For detailed accounts of Gale duality and its applications, we refer to the books of Grünbaum [17, Section 5.4], Ziegler [52, Chapter 6],  Matoušek [33, Section 5.6], McMullen and Shephard [36, Chapter 3], Björner et al. [6, Section 3.4], Pineda Villavicencio [40, Section 2.14], McMullen [35, Sections 6B, 6C] and to the papers of McMullen [34], Shephard [48], and Grünbaum and Shephard [18]. There are two versions of Gale duality: affine and linear. The version most commonly treated in the literature is the affine one; in the present work, however, we require linear Gale duality studied in [34, 48], which we now define.
@@
    if $A\in\mathbb{R}^{d^{\prime}\times n}$ is the matrix with columns $a_{1},\ldots,a_{n}$ and $B\in\mathbb{R}^{d^{\prime\prime}\times n}$ is the matrix with columns $b_{1},\ldots,b_{n}$, then every row of $A$ is orthogonal to every row of $B$. In matrix notation,
@@ 2.2. Gale coupling between two Gaussian projections
In this proof, we give an explicit construction of the coupling $(\mathbf{G},\mathbf{H})$. Let $X_{1},\ldots,X_{d^{\prime\prime}}$ be i.i.d. standard Gaussian points in $\mathbb{R}^{n}=\mathbb{R}^{d^{\prime}+d^{\prime\prime}}$. The linear Blaschke–Petkantschin formula [45, Theorem 7.2.1] states that for every nonnegative Borel function $h:(\mathbb{R}^{n})^{d^{\prime\prime}}\to[0,\infty)$,
@@
where $\nu_{n,d^{\prime\prime}}$ is the uniform probability distribution on $G(n,d^{\prime\prime})$, $\lambda_{L}$ is the Lebesgue measure on $L\in G(n,d^{\prime\prime})$, $\nabla_{d^{\prime\prime}}(z_{1},\ldots,z_{d^{\prime\prime}})$ denotes the $d^{\prime\prime}$-dimensional volume of the parallelepiped spanned by the vectors $z_{1},\ldots,z_{d^{\prime\prime}}$, and $B$ is a normalizing constant whose value can be found in [45, Eqn. (7.8), p. 271].
@@ 3.2. Expected face counts
As a simple consequence of Theorem 2.8, we recover the following result of Donoho and Tanner [11, Theorem 1.6]. Recall that Wendel’s formula [49] (see also [45, Theorem 8.2.1]) states that
@@
Taking up a suggestion of Gale, Schneider [43] proposed to study random polytopes $P$ obtained by choosing their (affine) Gale transform at random. (Throughout this remark, we use the term “polytope” to refer to its combinatorial type; only the combinatorial type is canonically determined by the Gale transform.) The model of Schneider [43, p. 643] depends on an even probability measure $\phi$; in the present remark we take $\phi$ to be the uniform distribution on the unit sphere in $\mathbb{R}^{k}$. We put $n=d+k$. Using the notation of Theorem 2.8, Schneider [43] considers a (d1)(d-1)-dimensional random polytope $P$ with $n$ vertices whose (affine) Gale transform is given by $\lambda_{1}V_{1},\ldots,\lambda_{n}V_{n}\in\mathbb{R}^{k}$ (with suitable $\lambda_{j}>0$) conditioned on the event $\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n})=\mathbb{R}^{k}$. By Example 2.4, this is the same conditioning event as $\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})\neq\mathbb{R}^{d}$. It then follows from Theorem 2.8 that the random cone $\mathcal{W}_{n,d}=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})$ conditioned on the event $\{\mathcal{W}_{n,d}\neq\mathbb{R}^{d}\}$, can be identified with the cone over $P$. Conversely, $P$ can be recovered as a bounded slice (cross-section) of this conditioned random cone. Such conditioned cones are called Cover–Efron cones in [20]. Consequently, the random polytopes studied in Schneider [43] have the same distribution as the bounded cross-sections of Cover–Efron cones considered in [20]. (As noted in [43, Remark on p. 648], these two objects in particular have the same expected face numbers.)
@@ 4.2. Moments of small order
In this section, we considered the cone $\mathcal{W}_{n,d}$ with $n=d$ only. Our methods do not allow to compute the variance of $\alpha(\mathcal{W}_{n,d})$ for $n>d$. What we can show is that for $n\geq d$,
(4.4) $\mathbb{E}[\alpha^{2}(\mathcal{W}_{n,d})]=\frac{\mathbb{E}[f_{1}^{2}(\mathcal{W}_{n+2,n+2-d})]-\mathbb{E}[f_{1}(\mathcal{W}_{n+2,n+2-d})]}{(n+2)(n+1)}.$
Proof of (4.4).
Let $k\in\mathbb{N}$. By Theorem 2.8, we may realize i.i.d. random vectors $U_{1},\ldots,U_{n+k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ together the i.i.d. random vectors $V_{1},\ldots,V_{n+k}\sim\mathrm{Unif}(\mathbb{S}^{n+k-d-1})$ on a common probability space such that the duality (2.6) holds. We have
$\mathbb{E}[\alpha^{k}(\mathcal{W}_{n,d})]$ $=\mathbb{P}\left[U_{n+1},\ldots,U_{n+k}\in-\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})\right]$
$=\mathbb{P}\left[\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n},U_{n+i})=\mathbb{R}^{d}\ \forall i=1,\ldots,k\right].$
Applying the duality (2.6), with $U_{i}$’s and $V_{j}$’s interchanged, to the right-hand side gives
$\mathbb{E}[\alpha^{k}(\mathcal{W}_{n,d})]=\mathbb{P}\left[\mathop{\mathrm{pos}}\nolimits\bigl(\{V_{n+1},\ldots,V_{n+k}\}\setminus\{V_{n+i}\}\bigr)\text{ is a face of }\operatorname{pos}(V_{1},\ldots,V_{n+k})\ \forall i=1,\ldots,k\right].$
If $k=2$, the condition on the right-hand side states that $\operatorname{pos}(V_{n+1})$ and $\operatorname{pos}(V_{n+2})$ are faces of $\operatorname{pos}(V_{1},\ldots,V_{n+2})$. So, using exchangeability,
$\mathbb{E}[\alpha^{2}(\mathcal{W}_{n,d})]$ $=\mathbb{P}\left[\mathop{\mathrm{pos}}\nolimits(V_{n+1}),\mathop{\mathrm{pos}}\nolimits(V_{n+2})\in\mathcal{F}_{1}\left(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2})\right)\right]$
$=\frac{1}{(n+2)(n+1)}\,\mathbb{E}\!\sum_{\begin{subarray}{c}i\neq j\\ i,j\in\{1,\ldots,n+2\}\end{subarray}}\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(V_{i})\in\mathcal{F}_{1}(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2}))\}}\,\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(V_{j})\in\mathcal{F}_{1}(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2}))\}}$
$=\frac{1}{(n+2)(n+1)}\,\mathbb{E}\!\left[f_{1}^{2}\left(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2})\right)-f_{1}\left(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2})\right)\right],$
and the proof is complete since $\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2})$ has the same law as $\mathcal{W}_{n+2,n+2-d}$. ∎
@@ 6.1. Statement of the limit theorem
Let us record a convex-dual analogue of Theorem 6.1. Let $U_{1},\ldots,U_{n}$ be independent random vectors, each uniformly distributed on $\mathbb{S}^{d-1}$, where $n>d$. The hyperplanes $U_{1}^{\perp},\ldots,U_{n}^{\perp}$ induce a conical tessellation of $\mathbb{R}^{d}$ into polyhedral cones. The number of cones is almost surely constant and equals $C(n,d)=2\sum_{\ell=0}^{d-1}\binom{n-1}{\ell}$, by the Steiner–Schläfli formula; see [45, Lemma 8.2.1]. Choose one of these cones uniformly at random and denote it by $\mathcal{S}_{n,d}$. The random cone $\mathcal{S}_{n,d}$ is called the random Schläfli cone; see [20, Definition 3.2].
@@ 6.2. Background from the theory of $U$-statistics
The proof of Theorem 6.1 relies on the theory of $U$-statistics. In this section, we introduce the necessary notation and recall several classical results on $U$-statistics; see the books of Lee [29], Korolyuk and Borovskikh [27], and Serfling [47, Chapter 5] for further background.
@@
For $n\geq m$, the variance of $\mathbb{U}_{n}$ is given by the exact formula (see [29, Chapter 1, § 1.3, Theorem 3] or [47, Section 5.2.1, Lemma A])
@@
Equivalently, $g_{r}(x_{1},\ldots,x_{r})=g_{0}$ for all $r\in\{1,\ldots,c-1\}$ ($\mu^{r}$-a.e.), but $g_{c}(X_{1},\ldots,X_{c})$ is not a.s. constant. If $h$ is degenerate of order $c$, Equation (6.7) implies (see, e.g., [47, Section 5.3.4])
@@
The next classical theorem characterizes the weak limit of the $U$-statisic $\mathbb{U}_{n}$ under degeneracy of order $2$; see [29, Section 3.2.2], [27, Chapter 4], [42, p. 168] or [47, Section 5.5.2, p. 194].
@@
The next lemma provides an explicit formula for the conditional expectation kernel $g_{c}(v_{1},\ldots,v_{c})$ of arbitrary order. The formula is expressed in terms of conic intrinsic volumes. To each closed convex cone $C\subseteq\mathbb{R}^{k}$ one can associate conic intrinsic volumes $\upsilon_{0}(C),\ldots,\upsilon_{k}(C)$, which are nonnegative and satisfy $\upsilon_{0}(C)+\cdots+\upsilon_{k}(C)=1$. We refer to [1, 2] and [44, Section 2.3] for their definition and properties.
@@
In the conditional probability, the only random element is the random matrix $\mathbb{O}$. To compute the conditional probability, we invoke the conic kinematic formula whose proof can be found in [1, Corollary 5.2] or in [44, Theorem  4.3.5]; for applications see also [1]. The formula states that for (deterministic) polyhedral cones $A,B\subseteq\mathbb{R}^{k}$ that are not both linear subspaces,
@@
Equations (6.13) and (6.14) are proved in [15, Lemma 5.1] (building on [20, Corollary 4.3]); where it is assumed that $n\geq d$. See also [24, Theorem 5.18] for an alternative approach. We sketch a proof of (6.15) for $n\in\{1,\dots,d\}$; for background on conic intrinsic volumes and hyperplane arrangements we refer to [44, Section 5.2]. For each sign vector $\varepsilon=(\varepsilon_{1},\dots,\varepsilon_{n})\in\{\pm 1\}^{n}$, consider the random cone $C_{\varepsilon}:=\mathop{\mathrm{pos}}\nolimits(\varepsilon_{1}U_{1},\dots,\varepsilon_{n}U_{n})$. By symmetry, $\mathbb{E}\,\upsilon_{k}(C_{\varepsilon})=\mathbb{E}\,\upsilon_{k}(\mathcal{W}_{n,d})$ for every $\varepsilon$. On the other hand, with probability $1$, the cones $C_{\varepsilon}$ are precisely the chambers of a hyperplane arrangement in $\mathop{\mathrm{lin}}\nolimits(U_{1},\dots,U_{n})$ that has the same intersection poset as the coordinate hyperplane arrangement in $\mathbb{R}^{n}$, whose chambers are the $2^{n}$ orthants. Since the $k$th conic intrinsic volume of an orthant in $\mathbb{R}^{n}$ equals $\frac{1}{2^{n}}\binom{n}{k}$, the Klivans–Swartz formula [44, Theorem 5.1.4] yields $\sum_{\varepsilon\in\{\pm 1\}^{n}}\upsilon_{k}(C_{\varepsilon})=\binom{n}{k}$ a.s. Taking expectations gives $2^{n}\,\mathbb{E}\,\upsilon_{k}(\mathcal{W}_{n,d})=\sum_{\varepsilon\in\{\pm 1\}^{n}}\mathbb{E}\,\upsilon_{k}(C_{\varepsilon})=\binom{n}{k}$, which is precisely (6.15). ∎
@@
Let $U$ be uniformly distributed on $\mathbb{S}^{k-1}$. For all $v,w\in\mathbb{S}^{k-1}$ it is easy to check (see, e.g., [51, Lemma 6.7 on p. 144]) that